# Time-Optimal Flight with Safety Constraints and Data-driven Dynamics

### v1

@@ Abstract
Time-optimal quadrotor flight is an extremely challenging problem due to the limited control authority encountered at the limit of handling. Model Predictive Contouring Control (MPCC) has emerged as a leading model-based approach for time optimization problems such as drone racing. However, the standard MPCC formulation used in quadrotor racing introduces the notion of the gates directly in the cost function, creating a multi-objective optimization that continuously trades off between maximizing progress and tracking the path accurately. This paper introduces three key components that enhance the MPCC approach for drone racing. First and foremost, we provide safety guarantees in the form of a constraint and terminal set. The safety set is designed as a spatial constraint which prevents gate collisions while allowing for time-optimization only in the cost function. Second, we augment the existing first principles dynamics with a residual term that captures complex aerodynamic effects and thrust forces learned directly from real world data. Third, we use Trust Region Bayesian Optimization (TuRBO), a state of the art global Bayesian Optimization algorithm, to tune the hyperparameters of the MPC controller given a sparse reward based on lap time minimization. The proposed approach achieves similar lap times to the best state-of-the-art RL and outperforms the best time-optimal controller while satisfying constraints. In both simulation and real-world, our approach consistently prevents gate crashes with 100% success rate, while pushing the quadrotor to its physical limit reaching speeds of more than 80km/h.
@@ I Introduction
The last decade has seen a remarkable growth in the number of quadrotor applications. Many missions are time-critical, such as search and rescue, aerial delivery, flying cars, or space exploration [1, 2, 3]. Among them, drone racing has emerged as a testbed for research on time-optimal flight. Quadrotors are carefully engineered to push the limits of speed and agility, making drone racing an ideal benchmark for aerial performance and robustness at high speeds [4, 5].
State-of-the-art approaches to autonomous drone racing can be categorized into learning-based and optimization-based methods. Within learning-based methods, Reinforcement Learning (RL) has arisen as an attractive alternative to conventional planning and control algorithms, outperforming world-champion pilots [6, 7]. Unlike optimal control, RL uses data sampled through millions of trial-and-error interactions with the environment to optimize a controller. RL can manage sparse objectives and unstructured observation spaces, providing substantial flexibility and versatility in the controller design.
Figure 1: Long exposure picture of our quadrotor flying at more than 80km/h in a physical racing track. The safety set is shown in translucent white, expanding in gate-free regions and constricting at gate passages. The safety set is a crucial element of our approach, thanks to which we can consistently prevent gate collisions and achieve 100% success rate in real-world experiments, even when operating at the drone’s physical limits.
@@
As shown in [7], RL can directly optimize a task-level objective, eliminating the need for explicit intermediate representations such as trajectories. Furthermore, RL can leverage domain randomization in simulation to cope with model uncertainty, allowing the discovery of more robust control responses. These advantages have facilitated numerous breakthroughs in pushing quadrotor systems to their operational limits through RL [7, 6]. However, despite the empirical success, RL controllers lack theoretical guarantees. Integrating safety considerations into machine learning frameworks, especially within RL, remains an area of ongoing research [8, 9]. The combination of optimization-based and learning-based architectures to enhance safety in RL is emerging as a prominent topic within the robotics community [10, 11, 12].
@@
Conversely, optimization-based architectures have tackled the problem from a different perspective. In [13], the authors propose a method to find time-optimal trajectories through a predefined set of waypoints and show how they outperform expert drone-racing pilots. However, these trajectories take several hours to calculate, rendering them impractical for replanning in the face of model mismatch and unknown disturbances (e.g., drone model, gate positions, aerodynamic effects, wind gusts). Addressing this challenge necessitates an algorithm capable of generating time-optimal trajectories in real time. In [14, 15], a Model Predictive Contouring Control (MPCC) method is introduced to track non-feasible paths in a time-optimal fashion. By solving the complex time-allocation problem online at every time step, MPCC selects the optimal states and inputs that maximize the progress along the designated path. This approach yields a controller that shows great adaptation against model mismatch and unknown disturbances.
Several modifications to the MPCC cost function are necessary to adapt the MPCC formulation to the drone racing task. First, the notion of gates is introduced in the objective function by parameterizing the contour weight as a function of the progress. This comes with adding extra hyperparameters, which i) introduces a multi-objective cost function that forces a sub-optimal trade-off between performance and safety, ii) does not generalize to different gate configurations, and iii) requires complex tuning due to the increased number of parameters. Various strategies have been devised to address the tuning problem for contouring controllers. For instance, [16] employs a policy search technique to explore the high-level parameter space of the cost function. Similarly, [17] explores the parameter space using local Bayesian Optimization (BO). Although these methods have resulted in control capabilities surpassing those of human pilots, the approximations in the cost function are still exceedingly intricate, suggesting the possibility for more robust and scalable alternatives. Such alternatives would ideally achieve the desired outcomes by imposing constraints that consistently prevent gate collisions, avoiding the need to handcraft complex, high-parametric cost functions.
@@
Previous works have aimed to tackle this issue by introducing a tunnel-shaped constraint set around the trajectory of the quadrotor [18, 19, 20]. In particular, [20] presents a tunnel-shaped constraint set by adapting the quadrotor’s dynamics into a local Frenet-Serret reference frame, which naturally gives rise to position constraints in the shape of a tunnel. The authors validate their approach in simulation and show that they can fly different tracks at high speeds within the confines of the tunnel.
@@
This paper introduces three key components that enhance the state-of-the-art MPCC formulation for drone racing. First, we provide safety guarantees in the form of a constraint and terminal set. The safety set is designed as a spatial constraint that prevents gate collisions in the form of a prismatic tunnel, similar to [20]. The terminal set consists of a periodic, feasible centerline. This combination provides guarantees of recursive feasibility and inherent robustness. Second, we augment the existing first principles dynamics with a residual term that captures complex aerodynamic effects and thrust forces learned directly from real-world data. Third, we use Trust-Region Bayesian Optimization (TuRBO), a state-of-the-art global Bayesian Optimization algorithm, to tune the hyperparameters of the MPC controller given a sparse reward based on lap-time minimization. In both simulation and real-world experiments, we illustrate how combining these elements improves the controller’s performance and robustness beyond the state-of-the-art MPCC. Moreover, the performance of our method aligns with that of the top-performing RL controller, with the added benefit of incorporating safety into our formulation.
@@ II-A High-speed quadrotor flight
The literature on high-speed, highly agile quadrotor flight is categorized into two primary approaches: model-based and learning-based. A comprehensive survey of the literature on this subject can be found in [5].
The model-based category originates from polynomial planning and classical control techniques. Traditionally, these methods focused on harnessing the differential flatness property of quadrotors and leverage the use of polynomial for planning [21, 22, 23, 24]. More recently, optimization-based methods have achieved planning of time-optimal trajectories using the quadrotor dynamics and numerical optimization [13].
@@
Nonetheless, the substantial computational demand of these methods typically necessitates offline trajectory computation, rendering these approaches impractical for real time applications. The leading model-based solution for minimum-time flight is Model Predictive Contouring Control (MPCC) [14], which synthesizes the trajectory planning and control tasks into one. A significant benefit of optimization-based methods is their ability to incorporate safety through state and input constraints. Notably, works like [25, 20, 18, 19] have investigated the application of positional constraints in the form of gates or tunnels to prevent collisions.
@@
On the other hand, a collection of agile flight learning-based methodologies for autonomous racing have emerged, which aim to replace the traditional planning and control layers with a neural network [26, 27, 28]. These purely data-driven control strategies, such as model-free reinforcement learning, strive to circumvent the limitations of model-based controller design by learning effective controllers directly from interactions with the environment. For instance, Hwangbo et al. [29] demonstrated the application of a neural network policy for guiding a quadrotor through waypoints and recovering from challenging initialization setups. Koch et al. [30] employed model-free RL for low-level attitude control and showed that a learned low-level controller trained with Proximal Policy Optimization (PPO) outperformed a fully tuned PID controller on almost every metric. Lambert et al. [31] implemented model-based RL to train a hovering controller. The family of learning-based approaches is rapidly advancing, fueled by recent breakthroughs in quadrotor simulation environments. These advancements have resulted in test environments that facilitate the training, assessment, and zero-shot transfer of control policies to the real world.
The state of the art on realistic simulations for quadrotors is [32], which introduces a hybrid aerodynamic quadrotor model which combines blade element momentum theory with learned aerodynamic representations from highly aggressive maneuvers.
@@
Song et al. [33] employed deep RL to generate near-time-optimal trajectories for autonomous drone racing and tracked the trajectories by an MPC controller. More recently, [6, 7] have demonstrated that policies trained with model-free RL can achieve super-human performance at drone racing. However, none of these learning-driven approaches include safety considerations into their designs.
II-B Automatic controller tuning
The classic approach [34] for controller tuning analytically finds the relationship between a performance metric, e.g. tracking error or trajectory completion, and optimizes the parameters with gradient-based optimization [35, 36, 37]. However, expressing the long term measure (in our case the laptime and gate passing metrics) as a function of the tuning parameters is impractical and generally intractable. There exist different methods for automatically tuning controllers [38, 39, 40, 41, 42, 43]. Another line of work proposes to iteratively estimate the optimization function, and use the estimate to find optimal parameters [44, 45, 46]. There have also been RL-rooted methods to find the right set of hyperparameters for a controller [16, 33]. Other methods are rooted in Bayesian Optimization, such as [17, 47].
@@ II-C Data-driven models for MPC
Learning-based MPC [48, 49, 50, 51, 52, 53, 54, 55, 56] utilize real-world data to refine dynamic models or learn an objective function tailored for MPC applications. These strategies typically focus on learning dynamics for tasks where deriving an analytical model of the robot or their operational environments poses significant challenges. This is particularly relevant for highly dynamic tasks, such as aggressive autonomous driving around a loose-surface track, expemplified in  [57]. Another promising line of research  [58] is the development of specialized solvers designed for use withing a learning-based MPC framework. This direction takes advantage of zero-order Sequential Quadratic Programming (SQP) techniques, potentially enhancing the efficiency and effectiveness of the control solutions for complex, dynamic tasks. This includes adapting to unpredictable elements and optimizing performance over a wide range of operational conditions.
@@ III Preliminaries
In this section we introduce the nominal quadrotor dynamics model that is later extended with learned dynamics. Additionally, we present the basics of the MPCC algorithm from [14].
@@ III-A Quadrotor Dynamics
In this section, we describe the nominal dynamics $\bm{f}(\bm{x},\bm{u})$ used throughout this paper, where $\bm{x}\in\mathbb{R}^{13}$ is the state of the quadrotor and $\bm{u}\in\mathbb{R}^{4}$ is the input to the system. The state of the quadrotor described from the inertial frame $I$ to the body frame $B$. Therefore, $\bm{x}=[\bm{p}_{IB},\bm{q}_{IB},\bm{v}_{IB},\bm{w}_{B}]^{T}$ where $\bm{p}_{IB}\in\mathbb{R}^{3}$ is the position, $\bm{q}_{IB}\in\mathbb{SO}(3)$ is the unit quaternion that describes the rotation of the platform, $\bm{v}_{IB}\in\mathbb{R}^{3}$ is the linear velocity vector, and $\bm{\omega}_{B}\in\mathbb{R}^{3}$ are the bodyrates in the body frame.
The input to the system is given as the collective thrust $\bm{f}_{T}=[0~{}~{}0~{}~{}f_{Bz}]^{T}$ and body torques $\bm{\tau}_{T}$. For ease of readability, we drop the frame indices, as they are consistent throughout the description. The nominal dynamic equations are
@@
where $\odot$ represents the Hamilton quaternion multiplication, $\mathbf{R}(\bm{q})$ the quaternion rotation, $m$ the quadrotor’s mass, and $\mathbf{J}$ the quadrotor’s inertia.
Additionally, the input space given by $\bm{f}$ and $\bm{\tau}$ is decomposed into single rotor thrusts $\bm{f}=[f_{1},f_{2},f_{3},f_{4}]$ where $f_{i}$ is the thrust at rotor $i\in\{1,2,3,4\}$.
$\bm{f}_{T}$ $=\begin{bmatrix}0\\ 0\\ \sum f_{i}\end{bmatrix}$ and $\text{and }\bm{\tau}_{T}$ $=\begin{bmatrix}l/\sqrt{2}(f_{1}+f_{2}-f_{3}-f_{4})\\ l/\sqrt{2}(-f_{1}+f_{2}+f_{3}-f_{4})\\ c_{\tau}(f_{1}-f_{2}+f_{3}-f_{4})\end{bmatrix},$ (2)
with the quadrotor’s arm length $l$ and the rotor’s torque constant $c_{\tau}$.
In addition to these nominal dynamics, we introduce the full dynamic model $\bm{\hat{f}}(\bm{x},\bm{u})=\bm{f}(\bm{x},\bm{u})+\bm{g}(\bm{x},\bm{u})$, which is augmented with a residual term $\bm{g}(\bm{x},\bm{u})$ that captures unmodeled terms such as aerodynamic effects and thrust forces. The residual term is inferred using real world data.

### v2

# MPCC++: Model Predictive Contouring Control for Time-Optimal Flight with Safety Constraints

@@ Abstract
Quadrotor flight is an extremely challenging problem due to the limited control authority encountered at the limit of handling. Model Predictive Contouring Control (MPCC) has emerged as a promising model-based approach for time optimization problems such as drone racing. However, the standard MPCC formulation used in quadrotor racing introduces the notion of the gates directly in the cost function, creating a multi-objective optimization that continuously trades off between maximizing progress and tracking the path accurately. This paper introduces three key components that enhance the state-of-the-art MPCC approach for drone racing. First and foremost, we provide safety guarantees in the form of a track constraint and terminal set. The track constraint is designed as a spatial constraint which prevents gate collisions while allowing for time optimization only in the cost function. Second, we augment the existing first principles dynamics with a residual term that captures complex aerodynamic effects and thrust forces learned directly from real-world data. Third, we use Trust Region Bayesian Optimization (TuRBO), a state-of-the-art global Bayesian Optimization algorithm, to tune the hyperparameters of the MPCC controller given a sparse reward based on lap time minimization. The proposed approach achieves similar lap times to the best-performing RL policy and outperforms the best model-based controller while satisfying constraints. In both simulation and real world, our approach consistently prevents gate crashes with 100% success rate, while pushing the quadrotor to its physical limits reaching speeds of more than 80km/h.
Supplementary Material
A narrated video with real-world experiments is available at:
https://youtu.be/sbKe9emghtM
@@ I Introduction
The last decade has seen a remarkable growth in the number of quadrotor applications. Many missions are time-critical, such as search and rescue, aerial delivery, flying cars, or space exploration [1, 2, 3]. Among them, drone racing has emerged as a testbed for research on time-optimal flight. Quadrotors are carefully engineered to push the limits of speed and agility, making drone racing an ideal benchmark for aerial performance and safety at high speeds [4, 5]. State-of-the-art approaches to autonomous drone racing can be categorized into learning-based and optimization-based methods.
@@
Within learning-based methods, Reinforcement Learning (RL) has arisen as an attractive alternative to conventional planning and control algorithms, outperforming world-champion pilots [6, 7]. Unlike optimal control, RL optimizes a controller using data sampled from numerous trial-and-error interactions with the environment. This approach can manage sparse objectives and unstructured observation spaces, providing substantial flexibility and versatility in the controller design. As shown in [7], RL can directly optimize a task-level objective, eliminating the need for explicit intermediate representations such as trajectories. Furthermore, it leverages domain randomization in simulation to cope with model uncertainty, allowing the discovery of more robust control responses. These advantages have facilitated numerous breakthroughs in pushing quadrotor systems to their operational limits. However, despite its empirical success, RL policies lack theoretical guarantees. Integrating safety considerations into learning-based frameworks remains an area of ongoing research [8, 9]. The combination of optimization-based and learning-based architectures to enhance safety is emerging as a prominent topic within the robotics community [10, 11, 12].
@@
Optimization-based architectures have tackled the problem from a different perspective. In [13], the authors propose a method to find time-optimal trajectories through a predefined set of waypoints and show how they outperform expert drone-racing pilots. However, these trajectories take several hours to calculate, rendering them impractical for replanning in the face of model mismatch and unknown disturbances (e.g., drone model, gate positions, aerodynamic effects, wind gusts). Addressing this challenge necessitates an algorithm capable of generating time-optimal trajectories in real-time. In [14, 15], a Model Predictive Contouring Control (MPCC) method is introduced to track non-feasible paths by maximizing progress along the designated path in a manner akin to time-optimal strategies. By solving the complex time-allocation problem online at every time step, MPCC selects the optimal states and inputs that enhance progress. This approach results in a controller that demonstrates great adaptatability in response to model mismatches and unknown disturbances.
Several modifications to the MPCC cost function are necessary to adapt the formulation to the drone racing task. In [14], the concept of gates is introduced in the cost function by parameterising the contour weight as a multivariate Gaussian function. Specifically, this adaptation employs a Gaussian function at each gate to scale the contour weight in the proximity of the gates. This ensures that the drone closely follows the reference path when navigating through the gate. However, this approach presents several shortcomings. First, employing a multivariate Gaussian to model the contour weight significantly increases the number of hyperparameters, making manual tuning a highly challenging task [16]. Secondly, while this contour function encourages the drone’s proximity to the designated path, it does not explicitly prevent gate collisions, resulting in a sub-optimal trade-off between performance and safety. Moreover, the contour function requires specific tuning for each gate configuration, which restricts its flexibility when gate positions change and increases the dimensionality of the parameter space with the number of gates.
Although this method has resulted in control capabilities surpassing those of human pilots, the approximations in the cost function are exceedingly intricate, suggesting the possibility for more robust and scalable alternatives. Such alternatives would ideally achieve the desired outcomes by imposing constraints that consistently prevent gate collisions, avoiding the need to handcraft complex, high-parametric cost functions.
@@
Previous works have aimed to tackle this issue by introducing a spatial constraint around the trajectory of the quadrotor [17, 18, 19]. In particular, [19] adapts the quadrotor’s dynamics to a local Frenet-Serret frame, which naturally gives rise to tunnel-like position constraints. However, this spatial reformulation of the dynamics requires using first and second derivatives of the coordinates in the Frenet-Serret frame, which introduces singularities that require careful handling.
@@
This paper introduces three key components that enhance the state-of-the-art MPCC formulation for drone racing [14]. First, we provide safety guarantees in the form of a track constraint and terminal set. The track constraint is designed as a spatial constraint that prevents gate collisions in the form of a prismatic tunnel. Unlike [19], our formulation retains the dynamics in their standard Euclidean form, avoiding the complex reformulation into the Frenet-Serret frame and its associated singularities. The terminal set consists of a periodic, feasible trajectory. This combination provides guarantees of recursive feasibility and inherent robustness. Second, we augment the existing first principles dynamics with a residual term that captures complex aerodynamic effects and thrust forces inferred directly from real-world data. Third, we show that using Trust-Region Bayesian Optimization (TuRBO) to tune the hyperparameters of the MPCC controller results in superior performance compared to previous work [16]. In both simulation and real-world experiments, we illustrate how combining these elements improves the controller’s performance and safety beyond the state-of-the-art MPCC. Moreover, the performance of our method aligns with that of the best-performing RL policy, with the added benefit of incorporating safety into our formulation.
@@ II-A High-speed quadrotor flight
The literature on high-speed quadrotor flight is categorized into two primary approaches: model-based and learning-based. A comprehensive survey of the literature on this subject can be found in [5].
The model-based category originates from polynomial planning and classical control techniques. Traditionally, these methods focused on harnessing the differential flatness property of quadrotors and leverage the use of polynomial for planning  [20, 21, 22, 23]. More recently, optimization-based methods have achieved planning of time-optimal trajectories using the quadrotor dynamics and numerical optimization [13]. Nonetheless, the substantial computational demand of these methods typically necessitates offline trajectory computation, rendering these approaches impractical for real-time applications. Within model-based approaches, the problem is typically framed either as time minimization or as progress maximization, depending on the optimization problem’s cost function. Time minimization formulations incorporate the time variable directly into the cost function but often require a spatial reformulation using the Frenet-Serret frame. This introduces several complexities, such as handling the inherent nonlinearities and singularities that arise from the coordinate transformations [24, 25, 26]. Conversely, contouring methods propose employing progress maximization along a path as a proxy for time minimization. This approach simplifies the control problem by allowing for more robust and efficient solutions. This distinction underlines that although progress maximization does not address the time-optimal problem directly, it offers a practical alternative with significant success. Consequently, MPCC has shown promising results in achieving minimum-time flight [14].
@@
A significant benefit of optimization-based methods is their ability to incorporate safety considerations through state and input constraints. Notably, works like  [17, 18, 19, 27] have investigated the application of positional constraints in the form of gates or tunnels to prevent collisions.
@@
On the other hand, a collection of learning-based methodologies for autonomous racing have emerged, which aim to replace the traditional planning and control layers with a neural network [28, 29, 30]. These purely data-driven control strategies, such as model-free RL, strive to circumvent the limitations of model-based controllers by learning effective policies directly from interactions with the environment. For instance, the authors in [31] employ a neural network policy for guiding a quadrotor through waypoints and recovering from challenging initialization setups. In [32], model-free RL is employed for low-level attitude control, showing that a learned low-level controller trained with Proximal Policy Optimization (PPO) outperformed a fully tuned PID controller. In [33], model-based RL is used to train a hovering controller. The family of learning-based approaches is rapidly advancing, fueled by recent breakthroughs in quadrotor simulation environments. These advancements have resulted in test environments that facilitate the training, assessment, and zero-shot transfer of control policies to the real world. The state-of-the-art on realistic simulations for quadrotors is [34], which introduces a hybrid aerodynamic quadrotor model that combines Blade Element Momentum Theory (BEM) with learned aerodynamic representations from highly aggressive maneuvers.
@@
The authors of [35] employed deep RL to generate near time-optimal trajectories for autonomous drone racing and tracked the trajectories by an MPC controller. More recently, [6, 7] have demonstrated that policies trained with model-free RL can achieve super-human performance. However, none of these learning-based approaches include safety considerations into their designs.
II-B Tunnel MPC
The concept of employing spatial constraints to systematically prevent collisions is well-established in car racing, where tracks are clearly defined at all points. However, drone racing introduces a unique challenge: while the positions of the gates are fixed, the space in between remains open for exploration. Recent advancements in this field have introduced the idea of using a tunnel around the trajectory as a spatial constraint [17, 18, 19]. This method involves a spatial reformulation of the drone’s dynamics into the Frenet-Serret frame, focusing on the transverse distances from the drone’s current position to the track’s centerline. Such a state definition, while providing a natural mechanism to enforce tunnel constraints, necessitates a conversion of Euclidean coordinates into terms dependent on these transverse distances. Although this approach elegantly integrates the spatial constraints within the drone’s control framework and has shown promising results in simulations [19], including high-speed navigation across various tracks, it also presents significant challenges due to the spatial reformulation required. Specifically, it introduces singularities that must be carefully addressed to maintain a smooth and reliable flight trajectory.
@@ II-C Data-driven models for MPC
Learning-based MPC [36, 37, 38, 39, 40, 41, 42, 43, 44] utilize real-world data to refine dynamic models or learn an objective function tailored for MPC applications. These strategies typically focus on learning dynamics for tasks where deriving an analytical model of the robot or their operational environments poses significant challenges. This is particularly relevant for highly dynamic tasks, such as aggressive autonomous driving around a loose-surface track, exemplified in [45].
Another promising line of research is the development of specialized solvers designed for use within a learning-based MPC framework [46]. This direction takes advantage of zero-order Sequential Quadratic Programming (SQP) techniques, potentially enhancing the efficiency of the control solutions for complex tasks. This includes adapting to unpredictable elements and optimizing performance over a wide range of operational conditions.
II-D Automatic controller tuning
The classic approach [47] for controller tuning analytically finds the relationship between a performance metric, e.g. tracking error or trajectory completion, and the controller parameters. It then optimizes the parameters with gradient-based methods [48, 49, 50]. However, expressing the long term performance metric, such as the lap time, as a function of the tuning parameters is impractical and generally intractable. There exist different methods for automatically tuning controllers [51, 52, 53, 54, 55, 56]. RL methods [35] utilize trial-and-error to adapt controller settings to complex system dynamics without the need for explicit modeling. Similarly, BO [57] leverages probabilistic models to navigate and explore the parameter space of black-box functions, effectively identifying optimal settings with minimal data. Furthermore, Weighted Maximum Likelihood Estimation (WML) [16] employs policy search techniques to iteratively refine parameter estimates from observed data, aiming to pinpoint the most effective parameters for the cost function. These methods provide a range of techniques for controller tuning, each with its own set of strengths and limitations.
@@ III Preliminaries
In this section we introduce the nominal quadrotor dynamics model and present the basics of the MPCC algorithm from [14].
@@ III-A Quadrotor Dynamics
In this section, we describe the nominal dynamics $\bm{f}(\bm{x},\bm{u})$ where $\bm{x}\in\mathbb{R}^{13}$ is the state of the quadrotor and $\bm{u}\in\mathbb{R}^{4}$ is the input to the system. The state of the quadrotor is given by $\bm{x}=[\bm{p}_{I},\bm{q}_{IB},\bm{v}_{I},\bm{w}_{B}]^{T}$, where $\bm{p}_{I}\in\mathbb{R}^{3}$ is the position, $\bm{q}_{IB}\in\mathbb{SO}(3)$ is the unit quaternion that describes the rotation from the body to the inertial frame, $\bm{v}_{I}\in\mathbb{R}^{3}$ is the linear velocity vector, and $\bm{\omega}_{B}\in\mathbb{R}^{3}$ are the bodyrates in the body frame $B$. For ease of readability, we drop the frame indices, as they remain consistent throughout the description. The nominal dynamic equations are given by:
@@
where $\odot$ represents the Hamilton quaternion multiplication, $\mathbf{R}(\bm{q})$ the quaternion rotation, $m$ the quadrotor’s mass, $\mathbf{J}$ the quadrotor’s inertia, $\bm{f}_{T}$ the collective thrust, and $\bm{\tau}_{T}$ the body torques. The input space, given by $\bm{f}_{T}$ and $\bm{\tau}_{T}$, is further decomposed into single rotor thrusts $\bm{f}=[f_{1},f_{2},f_{3},f_{4}]$ as follows:
$\bm{f}_{T}$ $=\begin{bmatrix}0\\ 0\\ \sum f_{i}\end{bmatrix}$ and $\text{and }\bm{\tau}_{T}$ $=\begin{bmatrix}l/\sqrt{2}(f_{1}+f_{2}-f_{3}-f_{4})\\ l/\sqrt{2}(-f_{1}+f_{2}+f_{3}-f_{4})\\ c_{\tau}(f_{1}-f_{2}+f_{3}-f_{4})\end{bmatrix}$ (2)
where $f_{i}$ is the i-th rotor’s thrust, $l$ the quadrotor’s arm length, and $c_{\tau}$ the rotor’s torque constant.
In addition to the nominal dynamics, we introduce the full dynamic model $\bm{\hat{f}}(\bm{x},\bm{u})=\bm{f}(\bm{x},\bm{u})+\bm{g}(\bm{x},\bm{u})$, which is augmented with a residual term $\bm{g}(\bm{x},\bm{u})$ that captures unmodeled terms such as aerodynamic effects. The residual term is inferred from real-world data.